OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (x^-2y^0)^2(4x^3y^-3)^3 ______________________________ (3x^0y^2)^2(x^-1y^3)^2

OpenStudy (anonymous):

first step would be to get rid of everything written to the power of 0, because \[b^0=1\]

OpenStudy (mathteacher1729):

Here is the original problem in \(\LaTeX\): \[\frac{(x^{-2}y^0)^2(4x^3y^{-3})^3}{(3x^0y^2)^2(x^{-1}y^3)^2}\]

OpenStudy (anonymous):

does ^+3y^-3 mean that the 3y term is raised to that power?

OpenStudy (anonymous):

listen to math teach use latex for comples exp.

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

so start with \[\frac{(x^{-2})^2(4x^3y^{-3})^3}{(y^2)^2(x^{-1}y^3)^2}\]

OpenStudy (anonymous):

no that is a mistake. need \[(3y^2)^2\] in the denominator sorry

OpenStudy (anonymous):

\[(256/9)x^7y^{-19}\]

OpenStudy (anonymous):

So 9 is the numerator and 256 is the denominator... right??

OpenStudy (anonymous):

\[(x^4)(64x^9y^-9)\] \[(9y^4)(x^-2y^6)\]

OpenStudy (anonymous):

\[64x ^{13}y^{-9}\] \[9y^{10}x^{-2}\]

OpenStudy (anonymous):

\[64x^{15}\] \[9y^{19}\] should be your final answer